OpenStudy (idealist10):

Find a curve through (-1, 3) orthogonal to every parabola of the form y=1+cx^2 that it intersects.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

orthogonal slope =-1/(slope of a line)

OpenStudy (idealist10):

So how do I do this problem?

OpenStudy (anonymous):

take the derivative of equation this will give the tangent line equation. the slope of orthogonal =-1/tangent slope

OpenStudy (idealist10):

Take the derivative of y=1+cx^2 with respect to x?

OpenStudy (anonymous):

y1-y2=m(x1-x2) (x2,y2)=(-1,3) y-3=m(x+1) you need m which is equal -1/m

OpenStudy (anonymous):

yes

OpenStudy (idealist10):

So I got dy/dx=2cx, now what?

OpenStudy (anonymous):

This is essentially an initial-value version of your previous orthogonal trajectory problems. Set up your ODE, then solve for \(y\) to find a family of solution curves, and finally plug in \(x=-1\) and \(y=3\) to determine the particular curve for this problem.

OpenStudy (idealist10):

OMG! It was so easy! I was over-thinking. Thank you for the hint.